OpenStudy (anonymous):

Jessica is selling books during the summer to earn money for college. She earns a commission on each sale but has to pay for her own expenses. After a month of driving from neighborhood to neighborhood and walking door-to-door, she figures out that her weekly earnings are approximately a linear function of the number of doors she knocks on. She writes the equation of the function like this: E(x) = 10x - 30, where x is the number of doors she knocks on during the week and E(x) is her earnings for the week in dollars.Which of the following are reasonable interpretations of the y-intercept

OpenStudy (anonymous):

A. Her expenses are $30 per week. B. She can earn $10 per week even if she does not knock on any doors. C. She will lose $10 per week if she does not knock on any doors. D. If she does not knock on any doors at all during the week, she will lose $30.

OpenStudy (danjs):

E(x) = 10*x-30 Earned = 10*(doors) + (some starting value for Earnings, or loss)

OpenStudy (danjs):

if she goes to 0 doors, the earnings are -30

OpenStudy (danjs):

expenses

OpenStudy (danjs):

the 10 is how much earned per door, or slope rate of change of earnings